OpenStudy (anonymous):

Radical equations help? URGENT 5. x^2/5 • x^2/9 6. 81^-3/4 7. (9+2√6)(6-√6) 8. i^12 9. √x-8 + 8 = x 10. 4x^5/6 11. 4√98 12. (7+i)-(6-2i) 13. A fraction has a denominator of 5 + 2i. What is multiplied with the fraction to rationalize the fraction? Thank you so much

OpenStudy (loser66):

any idea?

OpenStudy (anonymous):

@Loser66 huh? no i don't

OpenStudy (loser66):

do you know that \(\huge a^b * a^c = a^{b+c}\)?

OpenStudy (anonymous):

i do now haha

OpenStudy (loser66):

so, just apply \(\huge x^{\frac{2}{5}}*x^{\frac{2}{9}}\)

OpenStudy (anonymous):

\[x ^{4/13}?\]

OpenStudy (loser66):

2/5+2/9 is not 4/13. It's \(38/45\)

OpenStudy (anonymous):

...how

OpenStudy (loser66):

make them have the same denominator by multiply first fraction by 9 and second one by 5 , then add them.

OpenStudy (loser66):